# Around the "Fundamental Theorem of Algebra" (extended version)

### v2

@@ A Laurent polynomial in $n$ variables is called real if it takes real values on the compact subtorus
Indeed, for centrally symmetric compact convex sets $A,B$, the inclusion $A\subseteq B$ is equivalent to the inequality $h_{A}\leq h_{B}$ of their support functions. For $\xi\in{\mathbb{R}}^{n}$, the support function of ${\rm conv}(\Lambda)$ is
@@ 4.1. Exponential sums of one variable
We begin the proof of Theorem 4.1 with an auxiliary construction. Let $K\subset{\mathbb{C}}$ be a smooth compact curve. Assume first that $K$ does not intersect the zero set of an exponential sum $g$. Let ${\rm Arg}_{g,K}$ denote a branch of the multivalued function $\arg g(z)$ defined along $K$. The differential $d{\rm Arg}_{g,K}$ is independent of the chosen branch. We set
@@
Let $f$ be an exponential sum and let $K\subset{\mathbb{C}}$ be a smooth compact curve. Then there exists a constant $C>0$ such that
$|{\rm Arg}_{f}(z+K)|\leq C\qquad\text{for all }z\in{\mathbb{C}}.$
@@
Let $\Lambda$ be the spectrum of $f$, and let $E$ denote the vector space of exponential sums with spectrum $\Lambda$. Denote by $\mathbb{E}$ its projectivization; points of $\mathbb{E}$ correspond to exponential sums defined up to multiplication by a nonzero scalar. Translations by $z\in{\mathbb{C}}$ act on $E$:
$(w\cdot g)(z)=g(z+w),\qquad w\in{\mathbb{C}},$
@@
$\rho\colon{\mathbb{C}}\to{\rm Aut}(\mathbb{E}).$
Suppose, to the contrary, that no such constant $C$ exists. Then there is a sequence $\{z_{k}\}\subset{\mathbb{C}}$ such that $|{\rm Arg}_{f}(z_{k}+K)|\to\infty$. Passing to a subsequence, we may assume that $\rho(z_{k})(f)$ converges in $\mathbb{E}$ to a limit $f_{\infty}$. It would then follow that $|{\rm Arg}_{f_{\infty}}(K)|=\infty$, which is impossible. ∎
@@
Let $\Delta_{1},\dots,\Delta_{N}$ be the sides of the Newton polygon $\Delta={\rm conv}(\Lambda)$, and let $r_{1},\dots,r_{N}$ be the rays generated by the outward normal vectors to these sides. For each $j$, let $U_{j,R}$ denote the RR–neighborhood of $r_{j}$ in $\mathbb{C}$}. For $R$ sufficiently large, all zeros of $f$ are contained in $\bigcup_{j=1}^{N}U_{j,R}$. It therefore suffices to prove that
$N_{j}(f,r)=\frac{r}{2\pi}\,{\rm length}(\Delta_{j})+O(1),$ (4.2)
where $N_{j}(f,r)$ denotes the number of zeros of $f$ in $U_{j,R}\cap B_{r}$. Let $D_{j}$ be the boundary of the domain $U_{j,R}\cap B_{r}$. By the argument principle,
$N_{j}(f,r)=\frac{1}{2\pi}\,{\rm Arg}_{f}(D_{j}).$
Choose coordinates $x,y$ in $\mathbb{C}$} so that the ray $r_{j}$ coincides with the positive yy–axis. Then $D_{j}$ is the union of four curves: vertical segments $V_{-}(r)$ and $V_{+}(r)$ at $x=\pm R$, an arc $c_{l}$ of the circle of radius $R$ lying below the xx–axis, and an arc $c_{u}(r)$ of the circle of radius $r$ lying above the segment with endpoints $(-R,\sqrt{r^{2}-R^{2}})$ and $(R,\sqrt{r^{2}-R^{2}})$.
For $R$ sufficiently large, the increment of the argument of $f$ along $V_{-}(r)\cup V_{+}(r)$ equals $\sqrt{r^{2}-R^{2}}\,{\rm length}(\Delta_{j})$. The curve $c_{l}$ is independent of $r$, while the increment of the argument along $c_{u}(r)$ is bounded uniformly in $r$ by Lemma 4.1. This proves (4.2) and completes the proof. ∎
@@ 4.2.1. Definitions and examples
(See [21]). A more precise result based on the Fewnomial Theory is given in [22]) Assume that the spectra $\Lambda_{1},\ldots,\Lambda_{n}$ are contained in ${\rm Re\>}({{\mathbb{C}}^{n}}^{*})$. Then, for any domain $D\subset{\rm Im\>}({\mathbb{C}}^{n})$,

### v3

# Around the ‘‘Fundamental Theorem of Algebra’’

@@ A Laurent polynomial in $n$ variables is called real if it takes real values on the compact subtorus
Indeed, for compact convex sets $A,B$, the inclusion $A\subseteq B$ is equivalent to the inequality $h_{A}\leq h_{B}$ of their support functions. For $\xi\in{\mathbb{R}}^{n}$, the support function of ${\rm conv}(\Lambda)$ is
@@ 4.1. Exponential sums of one variable
We begin the proof of Theorem 4.1 with an auxiliary construction.
Let $K\subset{\mathbb{C}}$ be a smooth compact curve. Assume first that $K$ does not intersect the zero set of an exponential sum $g$. Let ${\rm Arg}_{g,K}$ denote a branch of the multivalued function $\arg g(z)$ defined along $K$. The differential $d{\rm Arg}_{g,K}$ is independent of the chosen branch. We set
@@
Let $A$ and $K$ be a compact set and a smooth compact curve in $\mathbb{C}$}, respectively. Then, for any fixed exponential sum $f$, there exists $C>0$ such that
(1) $\forall z\in{\mathbb{C}}\colon\,|{\rm Arg}_{f}(z+K)|\leq C$
(2) For the number of zeros $N(f,z+A)$ of the exponential sum $f$ in the set $z+A$ it is true that $\forall z\in{\mathbb{C}}\colon\,N(f,z+A)\leq C$.
@@
Let $\Lambda$ be the spectrum of $f$, and let $E$ denote the vector space of exponential sums with spectrum $\Lambda$. Let $\mathbb{E}=\mathbb{P}(E)$ be its projectivization; thus points of $\mathbb{E}$ correspond to exponential sums defined up to multiplication by a nonzero scalar.
Translations by $w\in{\mathbb{C}}$ act naturally on $E$ via
$(w\cdot g)(z)=g(z+w),$
@@
$\rho\colon{\mathbb{C}}\longrightarrow\mathrm{Aut}(\mathbb{E}).$
Assume, towards a contradiction, that no such constant $C$ exists. Then there is a sequence $\{z_{k}\}\subset{\mathbb{C}}$ such that at least one of the quantities
$|{\rm Arg}_{f}(z_{k}+K)|,\qquad N(f,z_{k}+A)$
tends to infinity as $k\to\infty$.
Passing to a subsequence if necessary, we may assume that $\rho(z_{k})(f)$ converges in $\mathbb{E}$ to a limit $f_{\infty}$. It would then follow that either $|{\rm Arg}_{f_{\infty}}(K)|=\infty$ or $N(f_{\infty},A)=\infty$, which is impossible. This contradiction completes the proof. ∎
@@
Let $\Delta_{1},\dots,\Delta_{N}$ be the sides of the Newton polygon $\Delta={\rm conv}(\Lambda)$, and let $r_{1},\dots,r_{N}$ be the rays generated by the outward normal vectors to these sides. For each $j$, denote by $U_{j,R}$ the $R$-neighborhood of $r_{j}$ in $\mathbb{C}$}.
For $R$ sufficiently large, all zeros of $f$ are contained in $\bigcup_{j=1}^{N}U_{j,R}$. It therefore suffices to prove that, for each $j$,
$N(f,U_{j,R}\cap B_{r})=\frac{r}{2\pi}\,{\rm length}(\Delta_{j})+O(1).$ (4.2)
Let $D_{j}$ denote the boundary of the domain $U_{j,R}\cap B_{r}$. By the argument principle,
$N(f,U_{j,R}\cap B_{r})=\frac{1}{2\pi}\,{\rm Arg}_{f}(D_{j}).$
Choose coordinates $(x,y)$ in $\mathbb{C}$} so that the ray $r_{j}$ coincides with the positive $y$-axis. Then $D_{j}$ is the union of four curves: the vertical segments $V_{-}(r)$ and $V_{+}(r)$ with endpoints (R,0),(R,r)(-R,0),(-R,r) and $(R,0),(R,r)$, respectively, an arc $c_{l}$ of the circle of radius $R$ centered at the origin lying below the $x$-axis, and an arc $c_{u}(r)$ of the circle of radius $r$ centered at the origin lying above the segment with endpoints $(-R,\sqrt{r^{2}-R^{2}})$ and $(R,\sqrt{r^{2}-R^{2}})$.
Fix $R$ sufficiently large. We estimate the increment of the argument of $f$ along each component of $D_{j}$.
(1) The increment of the argument of exponential sum $f$ along $V_{-}(r)\cup V_{+}(r)$ equals $2\sqrt{r^{2}-R^{2}}\,{\rm length}(\Delta_{j})$.
(2) The increment of the argument along $c_{l}$ is $O(1)$.
(3) By Lemma 4.1(1), the increment of the argument along the segment $V(r)$ joining the endpoints of the arc $c_{u}(r)$ is $O(1)$.
(4) Let $S(r)$ be the circular segment bounded by the arc $c_{u}(r)$ and its chord. As $r$ increases, the set $S(r)$, up to translation, is monotone decreasing. Hence, by Lemma 4.1(2), the number of zeros of $f$ in $S(r)$ is $O(1)$.
Combining (1)–(4), we obtain the estimate (4.2), completing the proof. ∎
@@ 4.2.1. Definitions and examples
(See [21]. A more precise result based on the Fewnomial Theory is given in [22]) Assume that the spectra $\Lambda_{1},\ldots,\Lambda_{n}$ are contained in ${\rm Re\>}({{\mathbb{C}}^{n}}^{*})$. Then, for any domain $D\subset{\rm Im\>}({\mathbb{C}}^{n})$,